For each function,
i) identify the values of the parameters a, h, and k
ii) identify the transformations
iii) use transformations to graph the function and check that it is
correct with a table of values or a graphing calculator

f(x)=(x+3)^2-2
1
Expert's answer
2020-12-01T09:10:52-0500

f(x)=(x+3)^2-2 = x^2 + 6x + 7.

a= 1.

h=-b/2a=-6/2=-3.

k=( -3)^2 + 6 x (-3) +7 = -2.

Parabola